OpenStudy (anonymous):

Find the center, vertices, and foci of the ellipse with equation x squared divided by 16 plus y squared divided by 25 = 1. Center: (0, 0); Vertices: (0, -5), (0, 5); Foci: (0, -3), (0, 3) Center: (0, 0); Vertices: (-5, 0), (5, 0); Foci: (0, -4), (0, 4) Center: (0, 0); Vertices: (0, -5), (0, 5); Foci: (-4, 0), (4, 0) Center: (0, 0); Vertices: (-5, 0), (5, 0); Foci: (-3, 0), (3, 0)

OpenStudy (anonymous):

@eliassaab

OpenStudy (anonymous):

@phi

OpenStudy (anonymous):

will give best answer & medal!

OpenStudy (anonymous):

@ganeshie8

OpenStudy (anonymous):

The first choice is correct. Let me explain why

OpenStudy (anonymous):

The foci are on the y-axis, since 25> 16. \[c^2=25-16=9\\ c=\pm 3 \] Foci (0,3),(0,03) Vartices (0,5),{0,-5)

OpenStudy (anonymous):

Corrected misprints from the post above Foci (0,3),(0,-3) Vertices (0,5),{0,-5)

OpenStudy (anonymous):

thank u!!!

OpenStudy (anonymous):

YW